OpenStudy (anonymous):

Arthur is organizing his collection of DVD'S he has 120 altogether The Ratio of comedy films to action to horror is 2:8:5 How many of each type of DVD'S Does he have Comedy.....DVD'S Action......DVD'S Horror.......DVD'S

OpenStudy (anonymous):

total number of dvd's = 120 total ratio = 2+ 8 + 5 = 15 so comedy' share = (2/15) * 120 = 16 actions share = (8/15) * 120 = 64 horrors share = 5/15 * 120 = 40